OpenStudy (anonymous):

Perform the indicated operation and, if possible, simplify. Please show all of your work. 2/9+(-3/7)

OpenStudy (anonymous):

Take the LCM of 9 and 7 first..

OpenStudy (anonymous):

+(-) = -

OpenStudy (anonymous):

ist multiply denominator of 1st term with numerator of 2nd term multiply denominator of 2nd term with numerator of 1st term

OpenStudy (anonymous):

\[\frac{2}{9} - \frac{3}{7}\] LCM of 9 and 7 is 63..

OpenStudy (anonymous):

next add the terms in numerators also use +(-) = - next mul the denominators

OpenStudy (anonymous):

cancel terms in num and denominator if any

OpenStudy (anonymous):

\[\frac{2 \times 7 - 3 \times 9}{63} = \frac{14 - 27}{63} = ??\]

OpenStudy (anonymous):

13 4/7

OpenStudy (anonymous):

I think you are getting negative numerator.. 14 - 27 = ??

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

\[\frac{-13}{63}\]

OpenStudy (anonymous):

So this is your answer..